OpenStudy (anonymous):

Find the equation of the parabola that passes through each set of points. (0,-7) (7,-14) (-3,-19)

OpenStudy (anonymous):

I got it to the point where I know that c=-7 The two equations I need to combine to get a or b are -7=49a+7b and -12=9a-3b but what do I multiply one of the equations by in order to be able to eliminate a or b?

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

Is that first equation correct? It seems like maybe it should be 49a + 7b + 14 = 0 or 49a + 7b = -14.

OpenStudy (anonymous):

ooh got it thanks:)

OpenStudy (anonymous):

When you are solving 2 equations with 2 unknowns, remember you can multiply both equations by different numbers if you need to in order to eliminate the "a" or "b".